An air bubble in a glass slab (R.I.=1.5)appears to be at 0.06m and 0.04m when viewed normally from the two opposite sides.The thickness in m,of the slab is

Grade:11

2 Answers

Arun
4 years ago
we can write  ( real-depth / apparent-depth ) = refractive index = 1.5
hence real depth is 1.5 times of apparent depth. If from one side, bubble is visible at 6 cm depth then the real depth is 9.
Similarly from other side if the bubble is visible at depth 4 cm, then real depth is 6 cm.
Hence the thickness of glass slab is 9+6 = 15 cm.
Khimraj
4 years ago
 

Refractive index of slab, μ=1.5

Total apparent length in the glass =0.06+0.04=0.10m

Thickness of the slab, t=μ× Apparent length in the glass =1.5×0.10=0.15m